# Algebraic perturbation theory: traversable wormholes and generalized entropy beyond subleading order

### v1

@@ 3.3 General change in weighting of expectation values
$\langle{\Psi}|\widehat{a}\,e^{\widetilde{H}_{L}}g^{-1}(\widetilde{H}_{L})|{\Psi}\rangle=\langle{\Psi}|\widehat{a}\,\exp\{\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}\}\,g^{-1}(\widetilde{H}_{L}))|{\Psi}\rangle$ (39)
$\approx\langle{\Psi}|\widehat{a}\underbrace{\left[1+\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}+\frac{1}{2}\left(\widetilde{H}_{L,0}^{2}+\delta\widetilde{H}_{L,0}^{2}+\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}\right)\right]}_{[\ldots]}g^{-1}(\widetilde{H}_{L}))|{\Psi}\rangle\leavevmode\nobreak\ ,$
where in the second line we have expanded to second order, and $\{\mathcal{O}_{1},\mathcal{O}_{2}\}$ is the anticommutator of operators $\mathcal{O}_{1}$, $\mathcal{O}_{2}$. In the underbrace, the second-order expansion in square brackets has been denoted []delimited-[][$\ldots$][ … ] for compactness in the sequel. Note that there is no need to expand $g^{-1}(\widetilde{H}_{L})$, because the state of the $L^{2}(\mathbb{R})$ factor above, $|{g^{1/2}}\rangle$, is the state corresponding to the perturbed TFD, so these will eventually cancel as in (32), cf. footnote 11.
To proceed, it is important to note the distinction between $H_{L,0}$ and $\widetilde{H}_{L,0}$: the former is the non-normalized modular charge, which formally acts on the (left factor of the) Hilbert space of the QFT. The latter is a suitably normalized version identified as the charge ${X_{0}=\widetilde{H}_{L,0}}$, which acts only on $L^{2}(\mathbb{R})$. Therefore, the tilded operators in (39) can be moved outside the TFD expectation value. Relatedly, note that the unitary $U$ that deforms the TFD acts only on the QFT Hilbert space, i.e.,
$|{\widetilde{\Psi}}\rangle=U|{\Psi_{0}}\rangle|{g^{1/2}}\rangle\leavevmode\nobreak\ ,$ (40)
so that the $L^{2}(\mathbb{R})$ factor is unaffected. However, the energy eigenstates (denoted $\lambda$ above) are affected, since $H_{L,0}$ – and therefore $X=\widetilde{H}_{L,0}$ – encodes information about the ADM mass, and hence knows about the deformation.
@@
where on the third line, we inserted another complete basis, whereupon we exploited the fact that $\langle{\lambda^{\prime}}|f(X)|{\lambda}\rangle=\delta(\lambda^{\prime}-\lambda)f(\lambda)$ since these are the eigenstates of $X$. The final weight ${\lambda}\right|{[\ldots]}\left|{\lambda}\right>\approx\left<{\lambda}\right|{Ue^{X_{0}}U^{\dagger}}\left|{\lambda}\right>$ is precisely the factor of $e^{\lambda}$ in (34), i.e.,
$e^{\lambda(\lambda_{0})}\approx\left<{\lambda}\right|{1+\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}+\frac{1}{2}\left(\widetilde{H}_{L,0}^{2}+\delta\widetilde{H}_{L,0}^{2}+\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}\right)}\left|{\lambda}\right>\leavevmode\nobreak\ .$ (42)
@@
However, the expansion (42) is not yet complete: until this point, we have implicitly absorbed the intrinsic temperature associated to the modular group into the modular charge $\widetilde{H}_{L}$, i.e., we have set $\beta=1$ (cf. (23); see for example Haag:1996hvx ; Connes:1994hv for background discussion on this point). Since $\mathfrak{A}\neq\mathfrak{A}_{0}$, we must allow for the possibility that $\beta\neq\beta_{0}$; physically, this is because the QES that acts as the bifurcation point for the boost operator (or the generator of modular flow more generally) may change, and the temperature is set by the associated horizon area. Therefore, we first restore the factors of $\beta$ in (42) arising from the fact that really, $e^{X}=e^{\beta\widetilde{H}_{L}}$:
$e^{\lambda(\lambda_{0})}\approx\left<{\lambda}\right|{1+\beta\widetilde{H}_{L,0}+\beta\delta\widetilde{H}_{L,0}+\frac{1}{2}\beta^{2}\left(\widetilde{H}_{L,0}^{2}+\delta\widetilde{H}_{L,0}^{2}+\{\widetilde{H}_{L,0},\delta\widetilde{H}_{L,0}\}\right)}\left|{\lambda}\right>\leavevmode\nobreak\ .$ (43)
@@
$\beta=\beta_{0}+\delta\beta_{0}\leavevmode\nobreak\ ,$ (44)
where $\delta\beta_{0}$ includes changes to all orders in the perturbation. In general, $\delta\beta_{0}\sim\delta\widetilde{H}_{L,0}$ in that both of them will be a series of perturbative corrections that, at least in the case of the double-trace deformation below, will be controlled by the same expansion parameter $h$. That is, we may write
@@
$\delta\beta_{0}\sim\delta\widetilde{H}_{L,0}\sim O(h)\leavevmode\nobreak\ ,$ (46)
@@
$+\frac{1}{2}\beta_{0}^{2}\,\delta\widetilde{H}_{L,0}^{2}+\delta\beta_{0}\left(\delta\widetilde{H}_{L,0}+\beta_{0}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\frac{1}{2}\delta\beta_{0}\widetilde{H}_{L,0}^{2}\right)|{\lambda}\rangle\leavevmode\nobreak\ ,$
where the terms on the first, second, and third line are $O(1)$, $O(h)$, and $O(h^{2})$, respectively. In particular, the ellipsis on the first line indicates the exponential series at $O(1)$, which we can re-sum to obtain
@@
$+\frac{1}{2}\beta_{0}^{2}\,\delta\widetilde{H}_{L,0}^{2}+\delta\beta_{0}\left(\delta\widetilde{H}_{L,0}+\beta_{0}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\frac{1}{2}\delta\beta_{0}\widetilde{H}_{L,0}^{2}\right)|{\lambda}\rangle\leavevmode\nobreak\ .$
@@ 3.4 Second-order perturbation theory for von Neumann entropy
We now wish to apply our machinery above to compute the change in the von Neumann entropy (27) under a perturbation of the type (38). Again, while we have endeavored to keep the present discussion as general as possible, we will shortly consider the specific case of a double-trace deformation of the TFD in the next section, so the reader may wish to keep that situation in mind if she finds the physical context helpful.
Concretely, denoting the density state in the type II algebras on the right-hand side before and after the deformation, respectively, by $\widehat{\rho}_{R,0}$ and $\widehat{\rho}_{R}$, we wish to compute
$\Delta S=S(\widehat{\rho}_{R})-S(\widehat{\rho}_{R,0})=-\mathrm{tr}\,\widehat{\rho}_{R}\ln\widehat{\rho}_{R}+\mathrm{tr}\,\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}\leavevmode\nobreak\ ,$ (49)
as mentioned at the beginning of subsec. 3.3. Note however that while $\widehat{\rho}_{R}=U\widehat{\rho}_{R,0}U^{\dagger}$, this is not a unitary operation with respect to the right algebra $\widehat{\mathfrak{A}}_{R}$, as explained in subsec. 3.1, so the change in entropy will generically be non-zero. Indeed, from the discussion leading up to (35), this becomes
$\Delta S=-\int\!\mathrm{d}\lambda_{0}\left(e^{\lambda(\lambda_{0})}-e^{\lambda_{0}}\right)\left<{\Psi_{0}}\right|{\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ .$ (50)
It thus remains only to compute the weight $e^{\lambda(\lambda_{0})}$, the meaning of which as an operator expectation value was given to second order in (48). Note that the leading exponential factor will cancel in the final expression for the change in entropy, since
$e^{\lambda_{0}}=\left<{\lambda}\right|{e^{\beta_{0}\widetilde{H}_{L,0}}}\left|{\lambda}\right>\leavevmode\nobreak\ .$ (51)
As we will now see, $\delta\widetilde{H}_{L,0}$ admits an expression as an infinite series of nested commutators. This follows from choosing a form for the unitary $U$ in (38),
$U=e^{-i\tau H}\qquad\mathrm{with}\qquad\tau\coloneqq t-t_{0}\leavevmode\nobreak\ ,$ (52)
@@
$=\exp\left\{\sum_{n=0}^{\infty}\frac{[(-i\tau H)^{n},\widetilde{H}_{L,0}]}{n!}\right\}\leavevmode\nobreak\ ,$
@@
$e^{A}Be^{-A}=\sum_{n=0}^{\infty}\frac{[(A)^{n},B]}{n!}\leavevmode\nobreak\ ,$ (54)
where $[(A)^{n},B]$ is the iterated commutator
$[(A)^{n},B]\coloneqq\underbrace{[A,\ldots[A,[A}_{\textrm{$n$ times}},B]]\ldots]\leavevmode\nobreak\ ,\qquad\quad[(A)^{0},B]\coloneqq B\leavevmode\nobreak\ .$ (55)
@@
$H=H_{0}+\delta H_{0}\leavevmode\nobreak\ ,$ (56)
where by construction, $[H_{0},\widetilde{H}_{L,0}]=0$. Practically, this amounts to replacing $H$ by $\delta H_{0}$ in the BCH expression above:
$Ue^{\widetilde{H}_{L,0}}U^{\dagger}=\exp\left\{\sum_{n=0}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\right\}=\exp\left\{\widetilde{H}_{L,0}+\sum_{n=1}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\right\}\leavevmode\nobreak\ ,$ (57)
@@
$\delta\widetilde{H}_{L,0}=\sum_{n=1}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\approx-i\tau[\delta H_{0},\widetilde{H}_{L,0}]-\frac{\tau^{2}}{2}[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]\leavevmode\nobreak\ ,$ (58)
where we used the fact that $[\delta H_{0},\widetilde{H}_{L,0}]\sim O(h)$, and in the second step we have expanded to second order in the strength of the perturbation $h$. Thus, inserting this into (48), the perturbative expression for the weight to this order is
@@
where we have taken $\delta\beta_{0}\approx\beta_{1}+\beta_{2}$ as per (45), and again $\{\mathcal{O}_{1},\mathcal{O}_{2}\}$ denotes the anticommutator. On the first line, everything except the leading exponential term is $O(h)$, while terms on the second and third line are all $O(h^{2})$. Terms carrying factors of $\tau$ depend on the deformation, while those without depend only on the expansion of the inverse temperature (of course, some terms depend on both). Substituting this back into (50), we find that order-by-order, the difference in the von Neumann entropy is
$\Delta S^{(0)}=$ $0\leavevmode\nobreak\ ,$ (60)
$\Delta S^{(1)}=$ $\int\!\mathrm{d}\lambda_{0}\langle{\lambda}|\,i\tau\beta_{0}[\delta H_{0},\widetilde{H}_{L,0}]+\frac{i\tau}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}$
$\qquad\quad\quad-\beta_{1}\widetilde{H}_{L,0}\left(1+\beta_{0}\widetilde{H}_{L,0}\right)|{\lambda}\rangle\left<{\Psi_{0}}\right|{\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ ,$
$\Delta S^{(2)}=$ $\int\!\mathrm{d}\lambda_{0}\langle{\lambda}|\frac{\tau^{2}}{2}\Big{(}\beta_{0}[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]+\frac{1}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]\}$
$\qquad\quad\quad+\beta_{0}^{2}[\delta H_{0},\widetilde{H}_{L,0}]^{2}\Big{)}+i\tau\left(\beta_{1}[\delta H_{0},\widetilde{H}_{L,0}]+\beta_{0}\beta_{1}\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}\right)$
$\qquad\quad\quad-\frac{1}{2}\beta_{1}^{2}\widetilde{H}_{L,0}^{2}-\beta_{2}\widetilde{H}_{L,0}\left(1+\beta_{0}\widetilde{H}_{L,0}\right)|{\lambda}\rangle\left<{\Psi_{0}}\right|{\widehat{\rho}_{R,0}\ln\widehat{\rho}_{R,0}}\left|{\Psi_{0}}\right>\leavevmode\nobreak\ ,$

### v2

@@ 3.2 General change in weighting of expectation values
$\langle{\Psi}|\widehat{a}\,e^{\widetilde{H}_{L}}g^{-1}(\widetilde{H}_{L})|{\Psi}\rangle=\langle{\Psi}|\widehat{a}\,\exp\{\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}\}\,g^{-1}(\widetilde{H}_{L})|{\Psi}\rangle$ (32)
$\approx\langle{\Psi}|\widehat{a}\underbrace{\left[1+\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}+\frac{1}{2}\left(\widetilde{H}_{L,0}^{2}+\delta\widetilde{H}_{L,0}^{2}+\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}\right)\right]}_{[\ldots]}g^{-1}(\widetilde{H}_{L})|{\Psi}\rangle\penalty 10000\ ,$
where in the second line we have expanded to second order, and $\{\mathcal{O}_{1},\mathcal{O}_{2}\}$ is the anticommutator of operators $\mathcal{O}_{1}$, $\mathcal{O}_{2}$. In the underbrace, the second-order expansion in square brackets has been denoted $\ldots$] for compactness in the sequel. Note that there is no need to expand $g^{-1}(\widetilde{H}_{L})$, because this will cancel in the computation of the trace, cf. footnote 12.
To proceed, it is important to note the distinction between $H_{L,0}$ and $\widetilde{H}_{L,0}$: the former is the non-normalized modular charge, which formally acts on the (left factor of the) Hilbert space of the QFT. The latter is a suitably normalized version identified as the charge $X_{0}=\widetilde{H}_{L,0}$}, which acts now on $L^{2}(\mathbb{R})$. Its renormalization renders both its expectation value and its fluctuations finite in the large $N$ limit, thus making it a valid observable whose quantum description is given by the $L^{2}(\mathbb{R})$ factor. Therefore, the tilded operators in (32) can be moved outside the TFD expectation value. Relatedly, note that the unitary $U$ that deforms the TFD acts only on the QFT Hilbert space, i.e.,
$|{\widetilde{\Psi}}\rangle=U|{\Psi_{0}}\rangle|{g^{1/2}}\rangle\penalty 10000\ ,$ (33)
so that the $L^{2}(\mathbb{R})$ factor is unaffected. However, the energy eigenstates (denoted $\lambda$ above) are affected, since $H_{L,0}$ – and therefore $X=\widetilde{H}_{L,0}$ – encodes information about the ADM mass, and hence knows about the deformation (which is also encoded in the Jacobian mentioned above).
@@
$=\int\!\mathrm{d}\lambda\,\cancel{\langle g^{1/2}|\lambda\rangle}\left<{\lambda}\right|{[\ldots]}\left|{\lambda}\right>\cancel{g^{-1}(\lambda)}\cancel{\langle\lambda|g^{1/2}\rangle}\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right$>
@@
where on the third line, we inserted another complete basis, whereupon we exploited the fact that $\langle{\lambda^{\prime}}|f(X)|{\lambda}\rangle=\delta(\lambda^{\prime}-\lambda)f(\lambda)$ since these are the eigenstates of $X$. The final weight $\left<{\lambda}\right|{[\ldots]}\left|{\lambda}\right>\approx\left<{\lambda}\right|{Ue^{X_{0}}U^{\dagger}}\left|{\lambda}\right$> is precisely the factor of $e^{\lambda}$ in (28), i.e.,
@@
However, the expansion (35) is not yet complete: until this point, we have implicitly absorbed the intrinsic temperature associated to the modular group into the modular charge $\widetilde{H}_{L}$, i.e., we have set $\beta=1$ (cf. (16); see for example Haag:1996hvx ; Connes:1994hv for background discussion on this point). Since $\mathfrak{A}\neq\mathfrak{A}_{0}$, we must allow for the possibility that $\beta\neq\beta_{0}$; physically, this is because the QES that acts as the bifurcation point for the boost operator (or the generator of modular flow more generally) may change, and the temperature is set by the associated horizon area. This effect is familiar from standard quantum statistical mechanics, where under a perturbation to the potential, the temperature itself is modified at leading order by the expectation value of the perturbing operator. In the remainder of this subsection, we systematically include the corrections at higher orders in the temperature.
To do so, we first restore the factors of $\beta$ in (35) arising from the fact that really, $e^{X}=e^{\beta\widetilde{H}_{L}}$:
[…]
@@
$\beta=\beta_{0}+\delta\beta_{0}\penalty 10000\ ,$ (37)
where $\delta\beta_{0}$ includes changes to all orders in the perturbation. In general, $\delta\beta_{0}\sim\delta\widetilde{H}_{L,0}$ in that both of them will be a series of perturbative corrections that, at least in the case of the double-trace deformation in our example below, will be controlled by the same expansion parameter $h$. That is, we may write
@@
$\delta\beta_{0}\sim\delta\widetilde{H}_{L,0}\sim O(h)\penalty 10000\ ,$ (39)
@@
$+\frac{1}{2}\beta_{0}^{2}\,\delta\widetilde{H}_{L,0}^{2}+\delta\beta_{0}\left(\delta\widetilde{H}_{L,0}+\beta_{0}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\frac{1}{2}\delta\beta_{0}\widetilde{H}_{L,0}^{2}\right)|{\lambda}\rangle\penalty 10000\ ,$
where the terms on the first line are $O(1)$, those on the second second are minimum $O(h)$, and those on the third are minimum $O(h^{2})$. In particular, the ellipsis on the first line indicates the exponential series at $O(1)$, which we can re-sum to obtain
@@
$+\frac{1}{2}\beta_{0}^{2}\,\delta\widetilde{H}_{L,0}^{2}+\delta\beta_{0}\left(\delta\widetilde{H}_{L,0}+\beta_{0}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\frac{1}{2}\delta\beta_{0}\widetilde{H}_{L,0}^{2}\right)|{\lambda}\rangle\penalty 10000\ .$
@@
Finally, we return to the Jacobian, $\tfrac{\mathrm{d}\lambda}{\mathrm{d}\lambda_{0}}$ in (28). To compute, this, we note that, as just discussed, $\lambda_{0}\coloneqq\left<{\lambda}\right|{\beta_{0}\widetilde{H}_{L,0}}\left|{\lambda}\right$>}, and
$\lambda\coloneqq\left<{\lambda}\right|{\beta\widetilde{H}_{L}}\left|{\lambda}\right$> $=\left<{\lambda}\right|{\big(\beta_{0}+\delta\beta_{0}\big)\big(\widetilde{H}_{L,0}+\delta\widetilde{H}_{L,0}\big)}\left|{\lambda}\right$> (42)
$=\left<{\lambda}\right|{\lambda_{0}\left(1+\frac{\delta\beta_{0}}{\beta_{0}}+\frac{\delta\widetilde{H}_{L,0}}{\widetilde{H}_{L,0}}\right)+\delta\beta_{0}\,\delta\widetilde{H}_{L,0}}\left|{\lambda}\right$>
and that the variations $\delta\beta_{0}$, $\delta\widetilde{H}_{L,0}$ depend only on the external perturbation to the system (parametrized by $h$) and not on $\lambda_{0}$. Thus, the derivative with respect to $\lambda_{0}$ simply picks-out the factor in parentheses as the Jacobian. Then, the entire prefactor in (28) is
$\frac{\mathrm{d}\lambda}{\mathrm{d}\lambda_{0}}e^{\lambda(\lambda_{0})}=$ $\,\langle{\lambda}|\left[1+\frac{\delta\beta_{0}}{\beta_{0}}+\frac{\delta\widetilde{H}_{L,0}}{\widetilde{H}_{L,0}}\right]\Bigg[e^{\beta_{0}\widetilde{H}_{L,0}}+\left(\beta_{0}\,\delta\widetilde{H}_{L,0}+\delta\beta_{0}\widetilde{H}_{L,0}\right)$ (43)
$\quad+\frac{1}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\delta\beta_{0}\,\beta_{0}\widetilde{H}_{L,0}^{2}+\frac{1}{2}\beta_{0}^{2}\,\delta\widetilde{H}_{L,0}^{2}$
$\quad+\delta\beta_{0}\left(\delta\widetilde{H}_{L,0}+\beta_{0}\{\widetilde{H}_{L,0},\,\delta\widetilde{H}_{L,0}\}+\frac{1}{2}\delta\beta_{0}\widetilde{H}_{L,0}^{2}\right)\Bigg]|{\lambda}\rangle\penalty 10000\ .$
Note that in these expressions, factors of $\widetilde{H}_{L,0}^{-1}$ are to be understood formally as the corresponding inverse operator acting from the left, and that in general $\widetilde{H}_{L,0}^{-1},\delta\widetilde{H}_{L,0}]\neq 0$.
@@ 3.3 Second-order perturbation theory for von Neumann entropy
We now wish to apply our machinery above to compute the von Neumann entropy following a perturbation of the type (31). Again, while we have endeavored to keep the present discussion as general as possible, we will shortly consider the specific case of a double-trace deformation of the TFD in the next section, so the reader may wish to keep that situation in mind if she finds the physical context helpful.
Concretely, denoting the density operator in the type II algebras on the right-hand side before and after the deformation, respectively, by $\widehat{\rho}_{R,0}$ and $\widehat{\rho}_{R}$, we wish to compute
$S(\widehat{\rho}_{R})=-\mathrm{tr}\,\widehat{\rho}_{R}\ln\widehat{\rho}_{R}\penalty 10000\ ,$ (44)
as mentioned at the beginning of subsec. 3.2. Note however that while $\widehat{\rho}_{R}=U\widehat{\rho}_{R,0}U^{\dagger}$, this is not a unitary operation with respect to the right algebra $\widehat{\mathfrak{A}}_{R}$, as explained in subsec. A, so the entropy in the deformed algebra will generically differ. Indeed, from the discussion leading up to (28), this becomes
[…]
It thus remains only to compute the Jacobian and the weight $e^{\lambda(\lambda_{0})}$, the meaning of which as an operator expectation value was given to second order in (43). Note that to leading order, the exponential factor is the same as that in the undeformed state, since
[…]
As alluded above, $\delta\widetilde{H}_{L,0}$ admits an expression as an infinite series of nested commutators. This follows from choosing a form for the unitary $U$ in (31),
$U=e^{-i\tau H}\qquad\mathrm{with}\qquad\tau\coloneqq t-t_{0}\penalty 10000\ ,$ (47)
@@
$=\exp\left\{\sum_{n=0}^{\infty}\frac{[(-i\tau H)^{n},\widetilde{H}_{L,0}]}{n!}\right\}\penalty 10000\ ,$
@@
$e^{A}Be^{-A}=\sum_{n=0}^{\infty}\frac{[(A)^{n},B]}{n!}\penalty 10000\ ,$ (49)
where $(A)^{n},B]$ is the iterated commutator
$(A)^{n},B]\coloneqq\underbrace{[A,\ldots[A,[A}_{\textrm{$n$ times}},B]]\ldots]\penalty 10000\ ,\qquad\quad[(A)^{0},B]\coloneqq B\penalty 10000\ .$ (50)
@@
$H=H_{0}+\delta H_{0}\penalty 10000\ ,$ (51)
where by construction, $H_{0},\widetilde{H}_{L,0}]=0$. Practically, this amounts to replacing $H$ by $\delta H_{0}$ in the BCH expression above:
$Ue^{\widetilde{H}_{L,0}}U^{\dagger}=\exp\left\{\sum_{n=0}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\right\}=\exp\left\{\widetilde{H}_{L,0}+\sum_{n=1}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\right\}\penalty 10000\ ,$ (52)
@@
$\delta\widetilde{H}_{L,0}=\sum_{n=1}^{\infty}\frac{[(-i\tau\delta H_{0})^{n},\widetilde{H}_{L,0}]}{n!}\approx-i\tau[\delta H_{0},\widetilde{H}_{L,0}]-\frac{\tau^{2}}{2}[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]\penalty 10000\ ,$ (53)
where we used the fact that $\delta H_{0},\widetilde{H}_{L,0}]\sim O(h)$, and in the second step we have expanded to second order in the strength of the perturbation $h$. Thus, inserting this into (43) and collecting terms, the perturbative expression for the weight to this order is
$\begin{aligned} \frac{\mathrm{d}\lambda}{\mathrm{d}\lambda_{0}}e^{\lambda(\lambda_{0})}\approx&\langle{\lambda}|e^{\beta_{0}\widetilde{H}_{L,0}}+\frac{\beta_{1}}{\beta_{0}}e^{\beta_{0}\widetilde{H}_{L,0}}+\beta_{1}\widetilde{H}_{L,0}\left(1+\beta_{0}\widetilde{H}_{L,0}\right)\\ &-i\tau\Bigg[\beta_{0}[\delta H_{0},\widetilde{H}_{L,0}]+\widetilde{H}_{L,0}^{-1}[\delta H_{0},\widetilde{H}_{L,0}]e^{\beta_{0}\widetilde{H}_{L,0}}+\frac{1}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}\Bigg]\\ &-i\tau\left(\beta_{1}\widetilde{H}_{L,0}^{-1}[\delta H_{0},\widetilde{H}_{L,0}]\widetilde{H}_{L,0}\left(1+\beta_{0}\widetilde{H}_{L,0}\right)+2\beta_{1}[\delta H_{0},\widetilde{H}_{L,0}]+\frac{3}{2}\beta_{0}\beta_{1}\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}\right)\\ &-\frac{\tau^{2}}{2}\Bigg[2\beta_{0}\left(\beta_{0}+\widetilde{H}_{L,0}^{-1}\right)[\delta H_{0},\widetilde{H}_{L,0}]^{2}+\beta_{0}^{2}\widetilde{H}_{L,0}^{-1}[\delta H_{0},\widetilde{H}_{L,0}]\{\widetilde{H}_{L,0},\,[\delta H_{0},\widetilde{H}_{L,0}]\}\\ &\quad\quad+\beta_{0}[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]+\widetilde{H}_{L,0}^{-1}[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]e^{\beta_{0}\widetilde{H}_{L,0}}+\frac{1}{2}\beta_{0}^{2}\{\widetilde{H}_{L,0},\,[\delta H_{0},[\delta H_{0},\widetilde{H}_{L,0}]]\}\Bigg]\\ &+\frac{\beta_{2}}{\beta_{0}}e^{\beta_{0}\widetilde{H}_{L,0}}+\frac{3}{2}\beta_{1}^{2}\widetilde{H}_{L,0}^{2}+\beta_{2}\widetilde{H}_{L,0}\left(1+\beta_{0}\widetilde{H}_{L,0}\right)+\frac{\beta_{1}^{2}}{\beta_{0}}\widetilde{H}_{L,0}|{\lambda}\rangle\penalty 10000\ .\end{aligned}$ (54)
where we have taken $\delta\beta_{0}\approx\beta_{1}+\beta_{2}$ as per (38), and again $\{\mathcal{O}_{1},\mathcal{O}_{2}\}$ denotes the anticommutator. On the first two lines, everything except the leading exponential term is $O(h)$, while all terms on the third through last lines are $O(h^{2})$. Terms carrying factors of $\tau$ depend on the deformation, while those without depend only on the expansion of the inverse temperature (of course, some terms depend on both).
We note in passing that if we are content to work abstractly in the perturbation, i.e., with $\delta\beta_{0}=\sum_{i=1}\beta_{i}$ as in (38) and similarly $\delta\widetilde{H}_{L,0}=\sum_{i=1}E_{i}$ (where the subscript denotes terms of $O(h^{i})$, and we have used $E_{i}$ to avoid confusion with the various other $H$’s appearing here), then we may easily obtain the perturbative expression for the weight to all orders in $h$:
$\frac{\mathrm{d}\lambda}{\mathrm{d}\lambda_{0}}e^{\lambda(\lambda_{0})}=\left(1+\sum_{k=1}^{\infty}\left(\beta_{k}^{\prime}+E_{k}^{\prime}\right)\right)\sum_{n=0}^{\infty}\frac{\lambda_{0}^{n}}{n!}\left[1+\sum_{k=1}^{\infty}\left(\beta_{k}^{\prime}+E_{k}^{\prime}+\frac{1}{\lambda_{0}}\sum_{\ell=1}^{k}\beta_{\ell}E_{k-\ell}\right)\right]^{n}\penalty 10000\ ,$ (55)
where
$\beta_{k}^{\prime}\coloneqq\frac{\beta_{k}}{\beta_{0}}\penalty 10000\ ,\quad\quad E_{k}^{\prime}\coloneqq\frac{E_{k}}{E_{0}}\penalty 10000\ .$ (56)
For concreteness however, we will continue to limit ourselves to second-order in the sequel.
Hence, substituting (54) back into (45), we find that order-by-order, the perturbed von Neumann entropy is